# Optimized Speculative Sampling for GPU Hardware Accelerators

### v1

@@ Abstract
In this work, we optimize speculative sampling for parallel hardware accelerators to improve sampling speed. We notice that substantial portions of the intermediate matrices necessary for speculative sampling can be computed concurrently. This allows us to distribute the workload across multiple GPU threads, enabling simultaneous operations on matrix segments within thread blocks. Additionally, we use fast on-chip memory to store intermediate results, thereby minimizing the frequency of slow read and write operations across different types of memory. This results in profiling time improvements ranging from 6% to 13% relative to the baseline implementation, without compromising accuracy. To further accelerate speculative sampling, probability distributions parameterized by softmax are approximated by sigmoid. This approximation approach results in significantly greater relative improvements in profiling time, ranging from 37% to 94%, with a slight decline in accuracy. We conduct extensive experiments on both automatic speech recognition and summarization tasks to validate the effectiveness of our optimization methods.
@@
Dominik Wagner1, Seanie Lee2, Ilja Baumann1, Philipp Seeberger1, Korbinian Riedhammer1, Tobias Bocklet1,3 1Technische Hochschule Nürnberg Georg Simon Ohm, 2KAIST, 3Intel Labs dominik.wagner@th-nuernberg.de lsnfamily02@kaist.ac.kr {ilja.baumann, philipp.seeberger, korbinian.riedhammer, tobias.bocklet}@th-nuernberg.de
@@ 1 Introduction
Large foundational speech and language models based on autoregressive Transformer Vaswani et al. (2017) models have demonstrated remarkable proficiency across a variety of downstream tasks Hsu et al. (2021); Radford et al. (2022); Touvron et al. (2023b); Achiam et al. (2024). These models frequently increase in size, consequently requiring more memory and computational resources. However, downstream applications, such as dialogue systems, have strict wall-clock constraints and are often required to generate long sequences Pope et al. (2023); Chi et al. (2023); Fischer et al. (2024). Due to the sequential token generation in autoregressive decoding, latency increases with both the length of the sequence and the size of the model, resulting in a significant barrier to widespread deployment. On many general-purpose GPU hardware accelerator architectures, the increasing size of models leads to more read and write operations between high-bandwidth memory (HBM) and on-chip shared memory (SRAM) at each decoding step, necessitating more memory bandwidth to meet latency constraints Pope et al. (2023); Dao et al. (2022); Dao (2024). Consequently, the speed of autoregressive decoding becomes primarily limited by the available memory bandwidth and not by the number of computations that need to be executed on the dedicated hardware Shazeer (2019).
@@
In many cases, tokens may be accurately generated by much smaller models that require fewer resources. Motivated by this hypothesis, speculative sampling has been developed to accelerate autoregressive sampling Stern et al. (2018); Xia et al. (2023); Leviathan et al. (2023); Chen et al. (2023a). Speculative sampling employs a small draft model to generate tokens, which are potential future outputs of a larger target model. These drafted tokens are then verified in parallel by the target model, and only tokens that meet the validation criteria are retained as final outputs to ensure generation accuracy. This approach has been shown to significantly reduce the frequency of time-consuming operations, thereby improving inference latency Leviathan et al. (2023); Chen et al. (2023a).
@@
In this paper, we focus on optimizing the validation part of speculative sampling to further increase the inference speed. Inspired by recent advances in accelerating computations in the attention mechanism Dao et al. (2022); Dao (2024), we explore two faster methods for assessing drafted tokens by leveraging the parallelism capabilities of modern GPUs and reducing the number of slow read and write operations across the memory hierarchy. We identify that a significant portion of the intermediate matrices required for the sampling process can be computed independently. Exploiting this observation, we distribute the workload across multiple GPU threads and simultaneously compute portions of the intermediate output matrices within thread blocks. Additionally, we store intermediate results in fast shared memory and minimize the number of slow read and write operations to and from high-bandwidth memory. This optimization method is faster than the non-optimized baseline implementation and exact with regard to the decoding outputs, i.e., it generates the same outputs as the non-optimized method.
@@
To further accelerate speculative sampling, we propose using sigmoid as a pointwise approximation to softmax (Bridle, 1989), which is used to parameterize distributions of target and draft models. Since sigmoid is applied to logits in element-wise fashion, it can be computed in parallel and fused with other sampling-related computations. This enables significant acceleration of the overall process, but results in a small accuracy decline due to the non-exact nature of the method.
@@
We evaluate our two optimized algorithms on automatic speech recognition (ASR) and summarization tasks, covering draft model sizes between 166M and 2B parameters and target model sizes between 244M and 13B parameters. The exact optimization method reduces profiling time between 6% and 13% relative to the baseline implementation without compromising accuracy. Moreover, the non-exact optimization method improves profiling time by 37% to 94%, albeit with a small reduction in accuracy. We summarize our main contributions as follows:
@@
  • Exploration of sigmoid as a pointwise approximation to softmax in an even faster but non-exact variant of speculative sampling.
@@ 2 Related work
Techniques such as quantization  Dettmers et al. (2022); Bondarenko et al. (2023); Stock et al. (2021); Nagel et al. (2021), pruning Voita et al. (2019); Lagunas et al. (2021); Gromov et al. (2024) and knowledge distillation Sun et al. (2019); Sanh et al. (2019); Jiao et al. (2020); Hsieh et al. (2023) have proven effective in reducing inference latency with minimal performance impact. However, these approaches often require architectural changes or custom training procedures. Efforts specifically targeting the reduction of memory bandwidth bottlenecks during decoding include methods like multi-query attention, which aims to optimize memory usage per attention layer Shazeer (2019), or FlashAttention (Dao et al., 2022; Dao, 2024), which aims to reduce the number of read/write operations between HBM and SRAM on GPUs. In Pope et al. (2023), improvements in large-scale inference latency have been achieved by partitioning models and workload across multiple accelerators combined with various low-level optimizations to improve communication efficiency between devices.
@@
Speculative sampling approaches can be broadly categorized by the way drafting and verification is performed Xia et al. (2024a). Drafting refers to the efficient prediction of multiple future tokens with a draft model and verification refers to the methods used to verify the token sequence with the target model. Some works use specialized draft models Xia et al. (2023); Zhou et al. (2024). Others employ an existing smaller model from the same series Chen et al. (2023a); Spector and Re (2023); Leviathan et al. (2023); Yang et al. (2024) or leverage the target model directly for drafting, e.g., by skipping intermediate layers Zhang et al. (2023), using special look-ahead tokens Monea et al. (2023), or additional modeling heads Stern et al. (2018); Cai et al. (2024); Zhang et al. (2024). Verification approaches first supported greedy decoding Stern et al. (2018); Xia et al. (2023); Zhang et al. (2023) and were subsequently extended to support other methods such as nucleus sampling Leviathan et al. (2023); Chen et al. (2023a). Recently, methods to verify multiple draft sequences in parallel have also been explored Miao et al. (2024); Cai et al. (2024); Spector and Re (2023).
@@
Several studies have experimented with ReLU as alternatives to softmax in the attention (Bai et al., 2023b; Shen et al., 2023; Hron et al., 2020; Hua et al., 2022; Li et al., 2022; Wortsman et al., 2023). They either require training new models from scratch or maintain the computational overhead of gathering information along the full sequence length. Other softmax-related optimizations are tailored to reduce the memory requirement during training by computing only smaller fractions of the full softmax output in the backward pass Lee and Lee (2023) or leverage word frequencies in the training data to speed up computation Grave et al. (2017). Shim et al. (2017) approximate softmax by computing only a fraction of the full input with singular value decomposition. Other approximation methods are specifically designed for custom hardware such as field-programmable gate arrays Chen et al. (2023b) and application specific integrated circuits Geng et al. (2018).
@@ 3.1 Preliminaries
First, given the prefix $(x_{1},\ldots,x_{i+c-1})$, $\gamma\in\mathbb{N}_{+}$ draft tokens are sequentially sampled with $M_{q}$: $x_{i+c}\sim q(x|x_{<i+c})$ for $c=1,\ldots,\gamma$. The draft tokens are then evaluated using the target model $M_{p}$, a process that can be performed in parallel. Each draft token $x_{i+c}\in\mathcal{V}$ is accepted if $r_{c}\leq\tau_{c}(x_{i+c})$ for $c=1,\ldots,\gamma$. The terms $r_{c}$ and $\tau_{c}(x_{i+c})$ are computed as follows:
@@ 3.2.1 Exact optimization
We notice that the intermediate elements needed for speculative sampling can be computed concurrently within thread blocks and are largely independent of other thread blocks. Specifically, we can compute $(\tau_{c}(x))_{x\in\mathcal{V}}$ and parts of Eq. 3 in parallel. The kernel is tiled Lam et al. (1991), such that each thread block computes a tile of $n$ elements from each matrix at once. Threads within a block jointly load segments of the probability matrices $p(x|x_{x<i+c})$ and $q(x|x_{<i+c})$, whose full dimensions are $B\times\gamma\times\lvert\mathcal{V}\rvert$, into SRAM, effectively distributing the overhead of load latency, as discussed in more detail in Ryoo et al. (2008). $B$ denotes the batch size used during speculative decoding.
@@
By reorganizing the computations as illustrated in Fig. 1, batches of $p({x}|{x}_{<i+c})$ and $q({x}|{x}_{<i+c})$ are loaded only once from HBM into SRAM. Moreover, most operations are coupled within the kernel and performed in parallel while using fast SRAM to store intermediate values. Only the results necessary to produce token acceptance decisions are written to HBM. Therefore, read and write operations across the memory hierarchy are minimized, enabling faster speculative sampling.
@@ 3.2.2 Approximated optimization
Figure 2: Overview of the computations within each thread block for sigmoid approximation. Each set of logits is scaled by a minimum constant $\alpha$ and a maximum constant $\beta$. Sigmoid activations $\sigma$ are then computed and stored in SRAM for each segment of draft and target logits. Subsequently, the intermediate values $\hat{f}_{k}$, $\hat{a}_{k}$, $\hat{b}_{k}$, and $\hat{\tau}_{c_{k}}$ are computed analogous to Fig. 1. The resulting outputs are then used to update $\hat{\tau}_{c}$, $\hat{a}$, and $\hat{b}$ in HBM.
To further accelerate speculative sampling, we use sigmoid to approximate $p(x|x_{<i+c})$ and $q(x|x_{i+c})$, which are parameterized by softmax in the baseline implementation and the exact method described in § 3.2.1. Instead of treating $p(x|x_{<i+c})$ and $q(x|x_{i+c})$ as precomputed inputs to the kernel, the sigmoid approximation is tightly coupled with the other operations in the speculative sampling process. This integration within the kernel substantially accelerates the overall sampling procedure.
@@
Let ${z}_{p}(x|x_{<i+c})$ be the logits of the target model $M_{p}$ given the prefix $(x_{1},\ldots,x_{i+c-1})$. Similarly, let ${z}_{q}(x|x_{<i+c})$ be the logits of the draft model $M_{q}$. First, we rescale the logits using predefined constant values $\alpha<0$ and $\beta>0$, and then apply sigmoid to these scaled logits to approximate $p(x|x_{i+c})$ and $q(x|x_{<i+c})$ as follows:
@@
where $\sigma(x)=1/(1+\exp(-x))$. Although all values of $\hat{p}(x|x_{<i+c})$ and $\hat{q}(x|x_{<i+c})$ are positive, they do not sum to 1 and thus do not constitute valid probability distributions. Nonetheless, we rely on these approximations with the hope that they are sufficiently accurate for guiding token acceptance decisions. Using the approximations $\hat{p}$ and $\hat{q}$, we accept the draft token $x_{i+c}$ sampled from $M_{q}$ if $r_{c}\leq\hat{\tau}_{c}(x_{i+c})$:
@@ 4.1 Experimental setup
LibriSpeech clean Small.EN Distil Small.EN 0.08 0.08 0.09 11.7% 71.9%
@@
LibriSpeech clean Large V2 Distil Large V2 0.07 0.07 0.15 10.8% 83.9%
@@
CNN/DM - Llama2 7B Sheared Llama 1.3B 0.30 0.30 0.26 05.7% 78.0%
@@
XSum - Llama2 7B Sheared Llama 1.3B 0.20 0.20 0.18 11.1% 83.6%
@@
We evaluate accuracy and inference speed of our optimized speculative sampling on ASR and single-document summarization. For ASR, we measure word error rates (WERs) on the test portions of three English datasets: CommonVoice 16 (CV16) Ardila et al. (2020), LibriSpeech Panayotov et al. (2015), and TED-LIUM Rousseau et al. (2012). For summarization, we use the test portions of Extreme Summarization (XSum) Narayan et al. (2018) and CNN/Daily Mail (CNN/DM) Nallapati et al. (2016) to evaluate the quality of summaries generated by language models with ROUGE-1 (Lin, 2004). For all tasks, we use the PyTorch (Paszke et al., 2019) profiling tool to obtain execution times for performing speculative sampling. We measure the execution time within the entire call stack of the speculative sampling function, including any nested function call (e.g., softmax). The profiling times are summed over all decoding steps and examples in a dataset, before the relative improvement is calculated.
@@
We set the batch size $B$ to 1 and employ the same heuristic used in the baseline speculative sampling implementation in Transformers library (Wolf et al., 2020), to set the number of draft tokens $\gamma$. Initially, $\gamma$ is set to 5 and increases by 2 if all speculative tokens sampled from the draft model are accepted; otherwise, it decreases by 1. We set the maximum sequence length to 256 tokens for ASR and 100 tokens for summarization. For ASR, using sigmoid approximation, $\alpha$ and $\beta$ are set to $-10^{3}$ and $10^{3}$, respectively. In summarization experiments, we use $\alpha=-10^{4}$ and $\beta=10^{4}$. We set $n=1024$, i.e., the maximum available threads per block on the NVIDIA A100 GPU.
@@
We employ Whisper (Radford et al., 2022) as the target model series for the ASR task. We use both the multilingual 1.55B parameter whisper-large-v2 version and the English-only 244M parameter whisper-small.en version. For the summarization task, we use Llama2 7B/13B (Touvron et al., 2023b), Qwen 7B (Bai et al., 2023a), and Gemma 7B (Mesnard et al., 2024).
@@
Following Leviathan et al. (2023); Chen et al. (2023a) and Zhou et al. (2024) we either use smaller models of the same series or distilled versions of the target model for drafting. The draft model family for the ASR task is Distil-Whisper (Gandhi et al., 2023). In particular, we use the 166M parameter small.en and the 756M parameter distil-large-v2 versions. The draft model for Llama2 is Sheared-LLaMA (Xia et al., 2024b), a version of Llama2 pruned to 1.3B parameters. The draft models corresponding to Qwen and Gemma are the 500M and 2B parameter versions of the same series.
@@
We use the implementation of speculative sampling provided by the Transformers library (Wolf et al., 2020) (v4.38.2) in conjunction with PyTorch (v2.2.2) as our baseline. All models are loaded in FP16 and executed on A100 GPUs with 80GB HBM using the same compute node running CUDA 12.3 and NVIDIA device driver version 545.

### v2

@@ Abstract
In this work, we optimize speculative sampling for parallel hardware accelerators to improve sampling speed. We notice that substantial portions of the intermediate matrices necessary for speculative sampling can be computed concurrently. This allows us to distribute the workload across multiple GPU threads, enabling simultaneous operations on matrix segments within thread blocks. This results in profiling time improvements ranging from 6% to 13% relative to the baseline implementation, without compromising accuracy. To further accelerate speculative sampling, probability distributions parameterized by softmax are approximated by sigmoid. This approximation approach results in significantly greater relative improvements in profiling time, ranging from 37% to 94%, with a minor decline in accuracy. We conduct extensive experiments on both automatic speech recognition and summarization tasks to validate the effectiveness of our optimization methods.
@@
Dominik Wagner1, Seanie Lee2, Ilja Baumann1, Philipp Seeberger1, Korbinian Riedhammer1, Tobias Bocklet1 1Technische Hochschule Nürnberg Georg Simon Ohm, 2KAIST dominik.wagner@th-nuernberg.de lsnfamily02@kaist.ac.kr {ilja.baumann, philipp.seeberger, korbinian.riedhammer, tobias.bocklet}@th-nuernberg.de
@@ 1 Introduction
Large foundational speech and language models based on autoregressive Transformer Vaswani et al. (2017) architectures have demonstrated remarkable proficiency across a variety of downstream tasks Hsu et al. (2021); Radford et al. (2022); Touvron et al. (2023b); Achiam et al. (2024). These models frequently increase in size, consequently requiring more memory and computational resources. However, downstream applications, such as dialogue systems, have strict wall-clock constraints and are often required to generate long sequences Pope et al. (2023); Chi et al. (2023); Fischer et al. (2024). Due to the sequential token generation in autoregressive decoding, latency increases with both the length of the sequence and the size of the model, resulting in a significant barrier to widespread deployment. On many general-purpose GPU hardware accelerator architectures, the increasing size of models leads to more read and write operations between high-bandwidth memory (HBM) and on-chip shared memory (SRAM) at each decoding step, necessitating more memory bandwidth to meet latency constraints Pope et al. (2023); Dao et al. (2022); Dao (2024). Consequently, the speed of autoregressive decoding becomes primarily limited by the available memory bandwidth and not by the number of computations that need to be executed on the dedicated hardware Shazeer (2019).
@@
In many cases, however, tokens may be accurately generated by much smaller models that require fewer resources. Motivated by this hypothesis, speculative sampling has been developed to accelerate autoregressive sampling Stern et al. (2018); Xia et al. (2023); Leviathan et al. (2023); Chen et al. (2023a). Speculative sampling employs a small draft model to generate tokens, which are potential future outputs of a larger target model. These drafted tokens are then verified in parallel by the target model, and only tokens that meet the validation criteria are retained as final outputs to ensure generation accuracy. This approach has been shown to significantly reduce the frequency of time-consuming operations, thereby improving inference latency Leviathan et al. (2023); Chen et al. (2023a).
@@
In this paper, we focus on optimizing the validation part of speculative sampling to further increase the inference speed. Inspired by recent advances in accelerating computations in the attention mechanism Dao et al. (2022); Dao (2024), we explore two faster methods for assessing drafted tokens by leveraging the parallelism capabilities of modern GPUs. We identify that a significant portion of the intermediate matrices required for the sampling process can be computed independently. Exploiting this observation, we distribute the workload across multiple GPU threads and simultaneously compute portions of the intermediate output matrices within thread blocks. This optimization method is faster than the non-optimized baseline implementation and exact with regard to the decoding outputs, i.e., it generates the same outputs as the non-optimized method.
@@
To further accelerate speculative sampling, we propose using sigmoid as an element-wise approximation to softmax (Bridle, 1989), which is used to parameterize distributions of target and draft models. Since sigmoid is applied to logits in element-wise fashion, it can be computed in parallel and fused with other sampling-related computations. This enables significant acceleration of the overall process, but results in a small accuracy decline due to the non-exact nature of the method.
@@
We evaluate our two optimized algorithms on automatic speech recognition (ASR) and summarization tasks, covering draft model sizes between 166M and 2B parameters and target model sizes between 244M and 13B parameters. The exact optimization method reduces profiling time between 6% and 13% relative to the baseline implementation without compromising accuracy. Moreover, the non-exact optimization method improves profiling time by 37% to 94%, albeit with a small reduction in accuracy. We summarize our main contributions as follows: [^1] The source code of our optimized sampling algorithm is available at https://github.com/dwgnr/optimized-speculative-sampling.
@@
  • Exploration of sigmoid as an element-wise approximation to softmax in an even faster but non-exact variant of speculative sampling.
@@ 2 Related work
Techniques such as quantization  Dettmers et al. (2022); Bondarenko et al. (2023); Stock et al. (2021); Nagel et al. (2021), pruning Voita et al. (2019); Lagunas et al. (2021); Gromov et al. (2024) and knowledge distillation Sun et al. (2019); Sanh et al. (2019); Jiao et al. (2020); Hsieh et al. (2023) have proven effective in reducing inference latency with minimal performance impact. However, these approaches often require architectural changes or custom training procedures. Efforts specifically targeting the reduction of memory bandwidth bottlenecks during decoding include methods like multi-query attention, which aims to optimize memory usage per attention layer Shazeer (2019), or FlashAttention (Dao et al., 2022; Dao, 2024), which aims to reduce the number of read/write operations between HBM and SRAM on GPUs. Pope et al. (2023), achieve improvements in large-scale inference latency by partitioning models and workload across multiple accelerators combined with various low-level optimizations to improve communication efficiency between devices.
@@
Speculative sampling approaches can be broadly categorized based on how drafting and verification are conducted Xia et al. (2024a). Drafting refers to the efficient prediction of multiple future tokens with a draft model and verification refers to the methods used to verify the token sequence with the target model. Some works use specialized draft models Xia et al. (2023); Zhou et al. (2024). Others employ an existing smaller model from the same series Chen et al. (2023a); Spector and Re (2023); Leviathan et al. (2023); Yang et al. (2024) or leverage the target model directly for drafting, e.g., by skipping intermediate layers Zhang et al. (2024b), using special look-ahead tokens Monea et al. (2023), or additional modeling heads Stern et al. (2018); Cai et al. (2024); Zhang et al. (2024a). Verification approaches first supported greedy decoding Stern et al. (2018); Xia et al. (2023); Zhang et al. (2024b) and were subsequently extended to support other methods such as nucleus sampling Leviathan et al. (2023); Chen et al. (2023a). Recently, methods to verify multiple draft sequences in parallel have also been explored Miao et al. (2024); Cai et al. (2024); Spector and Re (2023).
@@
Several studies have experimented with ReLU and sigmoid as alternatives to softmax in the attention mechanism (Bai et al., 2023b; Shen et al., 2023; Hron et al., 2020; Hua et al., 2022; Li et al., 2022; Wortsman et al., 2023; Ramapuram et al., 2024). They either require training new models from scratch or maintain the computational overhead of gathering information along the full sequence length. Other softmax-related optimizations are tailored to reduce the memory requirement during training by computing only smaller fractions of the full softmax output in the backward pass Lee and Lee (2023) or leverage word frequencies in the training data to speed up computation Grave et al. (2017). Shim et al. (2017) approximate softmax by computing only a fraction of the full input with singular value decomposition. Other approximation methods are specifically designed for custom hardware such as field-programmable gate arrays Chen et al. (2023b) and application specific integrated circuits Geng et al. (2018).
@@ 3.1 Preliminaries
First, given the prefix $(x_{1},\ldots,x_{i+c-1})$, $\gamma\in\mathbb{N}_{+}$ draft tokens are sequentially sampled with $M_{q}^{\texttt{draft}}$: $x_{i+c}\sim q(x|x_{<i+c})$ for $c=1,\ldots,\gamma$. The draft tokens are then evaluated using the target model $M_{p}^{\texttt{target}}$, a process that can be performed in parallel. Each draft token $x_{i+c}\in\mathcal{V}$ is accepted if $r_{c}\leq\tau_{c}(x_{i+c})$ for $c=1,\ldots,\gamma$. The terms $r_{c}$ and $\tau_{c}(x_{i+c})$ are computed as follows:
@@
$=\frac{a(x)}{b}$
We denote the numerator of Eq. 3 by $a(x)$ and the denominator by $b$, as these terms are treated separately in subsequent sections.
The underlying concept of speculative sampling is similar to rejection sampling (Neal, 2003). Intuitively, a new token for the target model $M_{p}^{\texttt{target}}$ is generated by first sampling from a smaller draft model $M_{q}^{\texttt{draft}}$, which shares the same support ($\mathcal{V}$) as $M_{p}^{\texttt{target}}$. The token sampled from $M_{q}^{\texttt{draft}}$ is then evaluated in parallel with $M_{p}^{\texttt{target}}$, and its acceptance is determined based on the probability ratio defined in Eq. 1. If the token is rejected, a new one is drawn using the modified distribution in Eq. 2.
@@ 3.2.1 Exact optimization
We notice that the intermediate elements needed for speculative sampling can be computed concurrently within thread blocks and are largely independent of other thread blocks. Specifically, we can compute $(\tau_{c}(x))_{x\in\mathcal{V}}$ and parts of Eq. 3 in parallel. The kernel is tiled Lam et al. (1991), such that each thread block computes a tile of $n$ elements from each matrix at once. Threads within a block jointly load segments of the probability matrices $p(x|x_{x<i+c})$ and $q(x|x_{<i+c})$, whose full dimensions are $B\times\gamma\times\lvert\mathcal{V}\rvert$, into SRAM, effectively distributing the overhead of load latency Ryoo et al. (2008). $B$ denotes the batch size used during speculative decoding.
@@
By reorganizing the computations as illustrated in Fig. 1, batches of $p({x}|{x}_{<i+c})$ and $q({x}|{x}_{<i+c})$ are loaded only once from HBM into SRAM. Moreover, most operations are coupled within the kernel and performed in parallel while using fast SRAM to store intermediate values. Only the results necessary to produce token acceptance decisions are written to HBM.
@@ 3.2.2 Approximated optimization
Figure 2: Overview of the computations within each thread block for sigmoid approximation. Each set of logits is scaled by a minimum constant $\alpha$ and a maximum constant $\beta$. Sigmoid activations $\sigma$ are then computed and stored in SRAM for each segment of draft and target logits. Subsequently, the intermediate values $\hat{f}_{k}(x)$, $\hat{a}_{k}(x)$, $\hat{b}_{k}$, and $\hat{\tau}_{c_{k}}(x)$ are computed analogous to Fig. 1. The resulting outputs are then used to update $\hat{\tau}_{c}(x)$, $\hat{a}(x)$, and $\hat{b}$ in HBM.
To further accelerate speculative sampling, we use sigmoid to approximate $p(x|x_{<i+c})$ and $q(x|x_{i+c})$, which are parameterized by softmax in the baseline implementation and the exact method described in Section 3.2.1. Instead of treating $p(x|x_{<i+c})$ and $q(x|x_{i+c})$ as precomputed inputs to the kernel, the sigmoid approximation is tightly coupled with the other operations in the speculative sampling process. This integration within the kernel substantially accelerates the overall sampling procedure.
@@
Let ${z}_{p}(x|x_{<i+c})$ be the logits of the target model $M_{p}^{\texttt{target}}$ given the prefix $(x_{1},\ldots,x_{i+c-1})$. Similarly, let ${z}_{q}(x|x_{<i+c})$ be the logits of the draft model $M_{q}^{\texttt{draft}}$. First, we rescale the logits using predefined constant values $\alpha<0$ and $\beta>0$, and then apply sigmoid to these scaled logits to approximate $p(x|x_{i+c})$ and $q(x|x_{<i+c})$ as follows:
@@
where $\sigma(x)=1/(1+\exp(-x))$. Although all values of $\hat{p}(x|x_{<i+c})$ and $\hat{q}(x|x_{<i+c})$ are positive, they do not sum to 1 and thus do not constitute valid probability distributions. Nonetheless, we rely on these approximations with the hope that they are sufficiently accurate for guiding token acceptance decisions. Using the approximations $\hat{p}$ and $\hat{q}$, we accept the draft token $x_{i+c}$ sampled from $M_{q}^{\texttt{draft}}$ if $r_{c}\leq\hat{\tau}_{c}(x_{i+c})$:
@@ 4.1 Experimental setup
LibriSpeech clean Whisper Small.EN Distil- Whisper Small.EN 0.08 0.08 0.09 11.7% 71.9%
@@
LibriSpeech clean Whisper Large V2 Distil- Whisper Large V2 0.07 0.07 0.15 10.8% 83.9%
@@
CNN/DM Llama2 7B Sheared Llama 1.3B 0.30 0.30 0.26 05.7% 78.0%
@@
Xsum Llama2 7B Sheared Llama 1.3B 0.20 0.20 0.18 11.1% 83.6%
@@
We evaluate accuracy and inference speed of our optimized speculative sampling on ASR and single-document summarization. For ASR, we measure word error rates (WERs) on the test portions of three English datasets: CommonVoice 16 (CV16) Ardila et al. (2020), LibriSpeech Panayotov et al. (2015), and TED-LIUM Rousseau et al. (2012). For summarization, we use the test portions of Extreme Summarization (Xsum) Narayan et al. (2018) and CNN/Daily Mail (CNN/DM) Nallapati et al. (2016) to evaluate the quality of summaries generated by language models with ROUGE-1 (Lin, 2004). Additional dataset details are provided in Section A.3.
For all tasks, we use the PyTorch (Paszke et al., 2019) profiling tool to obtain execution times for performing speculative sampling. We measure the execution time within the entire call stack of the speculative sampling function, including any nested function call (e.g. softmax). The profiling times are summed over all decoding steps and examples in a dataset, before the relative improvement is calculated.
@@
We set the batch size $B$ to 1 and employ the same heuristic used in the baseline speculative sampling implementation in the Transformers library (Wolf et al., 2020), to determine the number of draft tokens $\gamma$. Initially, $\gamma$ is set to 5 and increases by 2 if all speculative tokens sampled from the draft model are accepted; otherwise, it decreases by 1. We set the maximum sequence length to 256 tokens for ASR and 100 tokens for summarization. For ASR, using sigmoid approximation, $\alpha$ and $\beta$ are set to $-10^{3}$ and $10^{3}$, respectively. In summarization experiments, we use $\alpha=-10^{4}$ and $\beta=10^{4}$. We set $n=1024$, i.e., the maximum available threads per block on the NVIDIA A100 GPU.
@@
We employ Whisper (Radford et al., 2022) as the target model series for the ASR task. We use both the multilingual 1.55B parameter whisper-large-v2 version and the English-only 244M parameter whisper-small.en version of the model. For the summarization task, we use Llama2 7B/13B (Touvron et al., 2023b), Qwen 1.8B/7B (Bai et al., 2023a), and Gemma 7B (Mesnard et al., 2024). More details on the target models are provided in Section A.1.
@@
Following Leviathan et al. (2023), Chen et al. (2023a), and Zhou et al. (2024) we either use smaller models of the same series or distilled versions of the target model for drafting. The draft model family for the ASR task is Distil-Whisper (Gandhi et al., 2023). In particular, we use the 166M parameter small.en and the 756M parameter distil-large-v2 versions. The draft model for Llama2 is Sheared-LLaMA (Xia et al., 2024b), a version of Llama2 pruned to 1.3B parameters. The draft models corresponding to Qwen and Gemma are the 500M and 2B parameter versions of the same series. More details on the draft models are provided in Section A.2.
@@
We use the implementation of speculative sampling provided by the Transformers library (Wolf et al., 2020) (v4.38.2) in conjunction with PyTorch (v2.2.2) as our baseline. Unless stated otherwise, all models are loaded in FP16 and executed on A100 GPUs with 80GB HBM using the same compute node running CUDA 12.3 and NVIDIA device driver version 545.